# Optimal Rates in Continual Linear Regression via Increasing Regularization

### v1

@@ 1 Introduction
In continual learning, a learner encounters a sequence of tasks and aims to acquire new knowledge without “forgetting” what was learned in earlier tasks. Many algorithmic approaches have been proposed to address this challenge [see surveys in 44, 41]. However, a deeper theoretical understanding is still needed to clarify the principles governing continual learning and is essential for the practical and reliable deployment of such methods.
@@
  • We reduce continual linear regression with either explicit $\ell_{2}$ regularization or finite-step budget to incremental gradient descent on surrogate losses. These reductions enable unified analysis, even under arbitrary task orderings and non-realizable settings. Figure 1 visualizes our analytical flow.
@@
    • We introduce and analyze a generalized form of SGD for time-varying objectives and use it to propose an increasing regularization schedule that achieves the optimal rate of $\mathcal{O}(1/k)$, closing the existing gap between upper and lower bounds. See Table 1 for a summary.
@@
Notation: $k\!=\,$iterations; $d\!=\,$dimensions; $\bar{r},r_{\max}\!=\,$average/maximum data matrix ranks; ${a\wedge b\triangleq\min(a,b)}$.
@@
Fixed Ours (2025) $\frac{\log k}{k}$
Increasing Ours (2025) $\frac{1}{k}$
@@
Any Ours (2025) $\frac{1}{k}$ (**)
(*) They did not explicitly present such lower bounds, but the $M\!=\!2$ tasks construction from their proof of Theorem 10, can yield a $\Theta({1}/{k})$ random behavior by cloning those $2$ tasks $\left\lfloor{M/2}\right\rfloor$ times for any general $M$.
@@ 2 Setting: Continual linear regression with explicit or implicit regularization
We focus on the widely studied continual linear regression setting, which, despite its simplicity, often reveals key phenomena and interactions in continual learning [e.g., 10, 11, 30, 15, 35, 27, 47, 16].
@@
[…]
@@
The average—or population—loss is defined as the mean loss across all individual tasks $m\in M$. That is,
@@
Prior work analyzed not only the loss over all tasks but also the forgetting, or loss on seen tasks. Under the random orderings considered here, all of these quantities are typically close. We thus focus on average loss and discuss the others in Section 4.3.
@@
A large body of practical continual learning research focuses on mitigating forgetting by explicitly penalizing changes in parameter space [e.g., 24, 46, 2, 6]. Many employ regularization terms based on Fisher information [4], though others have found empirically that isotropic regularization often performs comparably well [31, 39]. Following recent theoretical work [e.g., 27, 12, 5, 28], we also focus on isotropic regularizers but discuss alternatives in Section 4.3.
@@
While the continual update step above admits a closed-form solution—useful for theoretical analysis [e.g., 27]—our paper does not directly leverage it. Instead, in Section 3, we reduce this step—which solves an entire task—to a single gradient step, thus enabling last-iterate SGD analysis of the scheme.
@@
The coefficients $\lambda_{t}$ and step counts $N_{t}$ in Schemes 1 and 2 control the “regularization strength” and how well the current loss is minimized. This is often seen as tuning the stability-plasticity tradeoff [17, 44]. Our paper identifies choices that lead to improved upper bounds.
@@ 3 Regularized continual linear regression reduces to Incremental GD
We define the Incremental Gradient Descent (IGD) scheme to cast both Schemes 1 and 2 within a unified framework, enabling a common analysis. The reductions and the flow in which we employ them are illustrated in Figure 1. At each iteration $t$, the algorithm performs a gradient step on the time-varying smooth convex function $f^{(t)}(\cdot;\tau_{t})$, selected by the ordering $\tau$, using step size $\eta_{t}$.
@@
We present two reductions that cast regularized and budgeted continual regression as special cases of incremental gradient descent. Proofs for this section are provided in Appendix B.
@@
For $t\in\left[{k}\right]$, define $f^{(t)}_{r},f^{(t)}_{b}$ as in Reductions 1 and 2, and recall the data radius $R\triangleq\max_{m\in\left[{M}\right]}\left\|{\mathbf{X}_{m}}\right\|_{2}$.
@@
    Finally, when the tasks are jointly realizable (see Assumption 4.1), the same ${\mathbf{w}}_{\star}$ minimizes all surrogate objectives simultaneously. That is,
@@ 4 Rates for realizable continual linear regression in random orderings
In this section, we focus on a setting in which all tasks can be perfectly solved by a single predictor—a common assumption [^2] Other theoretical works similarly assume an underlying linear model, but allow additive label noise. This, however, almost invariably requires assuming either i.i.d. features [15, 30, 3] or commutable covariance matrices across tasks [27, 28, 47]—whereas we allow arbitrary data matrices, enabling worst-case analysis. in theoretical continual learning [e.g., 11, 12, 25, 16, 22, 13]. This assumption simplifies analysis by allowing all iterates to be compared to a fixed predictor, ruling out task collections with inherent contradictions. Realizability often holds in highly overparameterized deep networks, which can typically be optimized to arbitrarily low loss. In the neural tangent kernel (NTK) regime [21, 7], such networks exhibit effectively linear dynamics that closely align with our analysis.
Assumption 4.1 (Joint realizability).
There exists an offline solution ${\mathbf{w}}_{\star}\in\mathbb{R}^{d}$ such that
@@
We study random orderings as a natural model of non-adversarial task sequences [37]. Such orderings avoid worst-case pathologies and allow reductions to standard stochastic tools. They are implicitly used when generating common random benchmarks (e.g., permuted or split datasets), and can also be induced algorithmically by random sampling. These settings have been studied empirically [26, 19] and theoretically [11, 12, 22, 13]. Table 1 compares known rates under random and cyclic orderings.
@@
Under random ordering with replacement, no algorithm can achieve a worst-case expected loss convergence rate faster than $\Omega(1/k)$. This result, which stems from the uncertainty over unseen tasks, is formally established in Theorem A.1 and serves as a baseline for evaluating the tightness of our upper bounds.
@@ 4.1 Near optimal rates via fixed, horizon-dependent regularization strength
We apply last-iterate convergence results for SGD to the surrogate losses used by IGD under random orderings. Specifically, using the results of Evron et al. [13] together with the smoothness and upper bound from Lemma 3.1, we establish:
@@
Assume a random with-replacement ordering over jointly realizable tasks. Then, for each of Schemes 1 and 2, the expected loss after $k\geq 1$ iterations is upper bounded as:
@@
    Fixed coefficient: For Scheme 1 with a regularization coefficient $\lambda>0$,
@@
    Fixed budget: For Scheme 2 with step size $\gamma\in(0,1/R^{2})$ and budget $N\in\mathbb{N}$,
@@
All proofs for this subsection are provided in App. C.
The rates established in Lemma 4.3 raise a natural question: What choice of the regularization strength—i.e., the regularization coefficient $\lambda$ or step count $N$—achieves the tightest bound?
@@
Assume a random with-replacement ordering over jointly realizable tasks. When the regularization strengths in Lemma 4.3 are set as follows:
@@
    Fixed coefficient: For Scheme 1, set regularization coefficient $\lambda\triangleq R^{2}(\ln k-1)$;
@@
    Fixed budget: For Scheme 2, choose step size $\gamma\in(0,1/R^{2})$ and set budget $N\triangleq\frac{\ln\left({1-\frac{1}{\ln k}}\right)}{\ln\left({1-\gamma R^{2}}\right)}$;
Then, under either Scheme 1 or Scheme 2, the expected loss after $k\geq 2$ iterations is bounded as:
@@
The rates in Lemmas 4.3 and 4.4 extend to random orderings without replacement; see App. C for details.
This marks a significant improvement over the $\mathcal{O}(1/k^{1/4})$ rate established by Evron et al. [13] for the unregularized scheme. By tuning the regularization strength, we gain control over the smoothness of the surrogate losses $f^{(t)}_{r}$ and $f^{(t)}_{b}$ in Reductions 1 and 2, allowing us to attain the $\mathcal{O}(\log k/k)$ rate that is optimal within the SGD framework used in their analysis. In contrast, their unregularized scheme lacked this flexibility, which made achieving such rates considerably more difficult and potentially out of reach. A similar rate can also be derived from the last-iterate bounds of Varre et al. [42], as the smoothness induced by our choice of regularization falls within the applicable regime of their results.
@@ 4.2 Optimal rates via increasing regularization
We present the first result in continual linear regression that achieves the optimal rate for the last iterate, matching the known lower bound. This is obtained by employing a schedule in which the regularization strength increases over time. We discuss these findings and their connections to prior work in Section 6. All proofs for this subsection are provided in App. D.
@@
Assume a random with-replacement ordering over jointly realizable tasks. Consider either Scheme 1 or Scheme 2 with the following time-dependent schedules:
@@
    Scheduled coefficient: For Scheme 1, set regularization coefficient $\lambda_{t}=\frac{13R^{2}}{3}\cdot\frac{k+1}{k-t+2}$;
@@
    For Scheme 2, choose step sizes $\gamma_{t}\in(0,1/R^{2})$ and set budget $N_{t}=\frac{3}{13\gamma_{t}R^{2}}\cdot\frac{k-t+2}{k+1}$;
Then, under either Scheme 1 or Scheme 2, the expected loss after $k\geq 2$ iterations is bounded as:
@@
Establishing the theorem requires a novel last-iterate bound in stochastic optimization, as no existing analysis yields a $\mathcal{O}(1/k)$ guarantee for last-iterate convergence in the realizable setting. A standard path to such rates is to use a decreasing step-size schedule. However, our setting is more nuanced: the quantities we control are the regularization strengths—i.e., the regularization coefficient or step budget in Scheme 1 or 2—which inherently modify the surrogate objectives $f^{(t)}_{r}$ and $f^{(t)}_{b}$ in Reductions 1 and 2.
To handle this, we analyze SGD applied to time-varying objectives—a generalization of standard SGD. For this analysis to yield meaningful guarantees, the evolving surrogates must closely approximate the original loss. Indeed, this condition holds, as verified by Lemma 3.1, thus enabling the application of the next lemma.
@@
Assume $\tau$ is a random with-replacement ordering over $M$ jointly-realizable convex and $\beta$-smooth loss functions $f(\cdot;m)\colon\mathbb{R}^{d}\to\mathbb{R}$. Define the average loss $f({\mathbf{w}})\triangleq\mathbb{E}_{m\sim\tau}f({\mathbf{w}};m)$. Let $k\geq 2$, and suppose $\left\{f^{(t)}(\cdot;m)\mid t\in[k],m\in[M]\right\}$ are time-varying surrogate losses that satisfy:
@@
Then, IGD (Scheme 3) with a diminishing step size that satisfies $\nu_{t}\leq\eta_{t}=\eta\left(\frac{k-t+2}{k+1}\right),~{}\forall t\in[k]$ for some $\eta\leq 3/(13\beta)$, guarantees the following expected loss bound:
@@ 4.3 Do not forget forgetting: Extension to seen-task loss
We now take the opportunity to briefly revisit our results through the lens of other quantities of interest beyond the average (training) loss defined in Definition 2.1.
@@
In theoretical work, however, it is essential to define such quantities explicitly. Doan et al. [10] defined forgetting at time $k$ as the drift in model outputs, e.g., $\frac{1}{k}\sum_{t=1}^{k}\left\|{\mathbf{X}_{\tau_{t}}({\mathbf{w}}_{k}-{\mathbf{w}}_{t})}\right\|^{2}$. Nevertheless, this can be large even if the model improves between times $t$ and $k$—that is, in the presence of positive backward transfer.
@@
However, in schemes like our regularized approaches (Schemes 1 and 2), where convergence is not achieved despite realizability, loss degradation can be negative due to backward transfer. As a result, it is sensitive to worst-case analytical “manipulations” and difficult to analyze theoretically.
@@
In App. D, we extend Theorem 4.6 from the average loss to the seen-task loss. Specifically, we show that increasing regularization also achieves an $\mathcal{O}(1/k)$ rate for the expected seen-task loss. But, is this the optimal rate for seen-task loss?
The next lemma shows that, at least under explicit isotropic regularization (Scheme 1), it is optimal. Proof in App. A. More precisely, under random task orderings, no regularization schedule yields a rate faster than $\mathcal{O}(1/k)$ for the expected seen-task loss. In Section 6, we discuss how non-isotropic regularization—at the cost of additional space complexity—can ensure a seen-task loss of zero.
Lemma 4.9 (Lower bound for seen-task loss under Scheme 1).
For any $d\geq 2$, initialization ${\mathbf{w}}_{0}\in\mathbb{R}^{d}$, and regularization coefficient sequence $\lambda_{1},\ldots,\lambda_{k}\geq 0$, there exists a set of jointly realizable linear regression tasks $\left\{\left({\mathbf{X}_{m},\mathbf{y}_{m}}\right)\right\}_{m=1}^{M}$ such that, under a with-replacement random task ordering, Scheme 1 incurs seen-task loss $\mathcal{L}_{1:k}({\mathbf{w}}_{k})=\Omega(1/k)$ with probability at least $1/10$.

### v2

@@ 1 Introduction
In continual learning, a learner encounters a sequence of tasks and aims to acquire new knowledge without “forgetting” what was learned in earlier tasks. Many algorithmic approaches have been proposed to address this challenge [see surveys in 45, 42]. However, a deeper theoretical understanding is still needed to clarify the principles governing continual learning and is essential for the practical and reliable deployment of such methods.
@@
  • We reduce continual linear regression with either explicit $\ell_{2}$ regularization or finite-step budget to incremental gradient descent on surrogate losses. These reductions enable unified analysis, even under arbitrary task orderings and non-realizable settings. Figure˜1 visualizes our analytical flow.
@@
    • We introduce and analyze a generalized form of SGD for time-varying objectives and use it to propose an increasing regularization schedule that achieves the optimal rate of $\mathcal{O}(1/k)$, closing the existing gap between upper and lower bounds. See Table˜1 for a summary.
@@
Notation: k=k\!=\,iterations; d=d\!=\,dimensions; $\bar{r},r_{\max}\!=\,$average/maximum data matrix ranks; $a\wedge b\triangleq\min(a,b)$}.
@@
Fixed Ours $\frac{\log k}{k}$
Increasing Ours $\frac{1}{k}$
@@
Any Ours $\frac{1}{k}$ (**)
(*) They did not explicitly present such lower bounds, but the 22-task construction from their proof of Theorem 10, can yield a $\Theta({1}/{k})$ random behavior by cloning those 22 tasks $\left\lfloor{M/2}\right\rfloor$ times for any general $M$.
@@ 2 Setting: Continual linear regression with explicit or implicit regularization
We focus on the widely studied continual linear regression setting, which, despite its simplicity, often reveals key phenomena and interactions in continual learning [e.g., 11, 12, 31, 16, 36, 28, 48, 17].
@@
Bold symbols are reserved for matrices and vectors. Denote the Euclidean (vectors) or spectral (matrices) norm by $\left\|{\cdot}\right\|$, and the Moore-Penrose inverse by $\mathbf{X}^{+}$. Finally, denote $\left[{n}\right]=1,.\thinspace\!.\thinspace\!.,n$.
@@
The average—or population—loss is defined as the mean training loss across all individual tasks $m\in M$. That is,
@@
Prior work analyzed not only the loss over all tasks but also the forgetting, or loss on seen tasks. Under the random orderings considered here, all of these quantities are typically close. We thus focus on average loss and discuss the others in Section˜4.3.
@@
A large body of practical continual learning research focuses on mitigating forgetting by explicitly penalizing changes in parameter space [e.g., 25, 47, 2, 6]. Many employ regularization terms based on Fisher information [4], though others have found empirically that $\ell_{2}$ (isotropic) regularization often performs comparably well [32, 40]. Following recent theoretical work [e.g., 28, 13, 5], we also focus on isotropic regularizers but discuss alternatives in Section 4.3.
@@
While the continual update step above admits a closed-form solution—useful for theoretical analysis [e.g., 28]—our paper does not directly leverage it. Instead, in Section˜3, we reduce this step—which solves an entire task—to a single gradient step, thus enabling last-iterate SGD analysis of the scheme.
@@
The coefficients $\lambda_{t}$ and step counts $N_{t}$ in Schemes˜1 and 2 control the “regularization strength” and determine how well the current loss is minimized—often seen as tuning the stability-plasticity trade-off [18, 45]. Our paper identifies choices that yield improved bounds.
@@ 3 Regularized continual linear regression reduces to Incremental GD
We define the Incremental Gradient Descent (IGD) scheme to cast both Schemes˜1 and 2 within a unified framework, enabling a common analysis. The reductions and the flow in which we employ them are illustrated in Figure˜1. At each iteration $t$, the algorithm performs a gradient step on the time-varying smooth convex function $f^{(t)}(\cdot;\tau_{t})$, selected by the ordering $\tau$, using step size $\eta_{t}$.
@@
We present two reductions that cast regularized and budgeted continual regression as special cases of incremental gradient descent. Proofs for this section are provided in Appendix C.
@@
For $t\in\left[{k}\right]$, define $f^{(t)}_{r},f^{(t)}_{b}$ as in Reductions˜1 and 2, and recall the data radius $R\triangleq\max_{m\in\left[{M}\right]}\left\|{\mathbf{X}_{m}}\right\|_{2}$.
@@
    Finally, when the tasks are jointly realizable (see Assumption˜4.1), the same $\mathbf{w}$$_{\star}$ minimizes all surrogate objectives simultaneously. That is,
@@ 4 Rates for realizable continual linear regression in random orderings
In this section, we focus on a setting in which the training data of all tasks can be perfectly fit by a single predictor—a common assumption [^2] Other theoretical works similarly assume an underlying linear model, but allow additive label noise. This, however, almost invariably requires assuming either i.i.d. features [16, 31, 3] or commutable covariance matrices across tasks [28, 29, 48]—whereas we allow arbitrary data matrices, enabling worst-case analysis. in theoretical continual learning [e.g., 12, 13, 26, 17, 23, 14]. This assumption simplifies analysis by allowing all iterates to be compared to a fixed predictor, ruling out task collections with inherent contradictions. Such realizability often holds in highly overparameterized deep networks, which can typically be optimized to arbitrarily low training loss. In the neural tangent kernel (NTK) regime [22, 7], such networks exhibit effectively linear dynamics that closely align with our analysis.
Assumption 4.1 (Joint realizability of training data).
There exists an offline solution ${\mathbf{w}}_{\star}\in\mathbb{R}^{d}$ s.t.,
@@
We note in passing that even under this assumption, models still suffer from forgetting prior expertise—sometimes catastrophically—as thoroughly discussed in Evron et al. [12].
@@
We study random orderings as a natural model of non-adversarial task sequences [38]. Such orderings avoid worst-case pathologies and allow reductions to standard stochastic tools. They are implicitly used when generating common random benchmarks (e.g., permuted or split datasets), and can also be induced algorithmically by random sampling. These settings have been studied empirically [27, 20] and theoretically [12, 13, 23, 14]. Table˜1 compares known rates under random and cyclic orderings.
@@
Under random ordering with replacement, no algorithm can achieve a worst-case expected loss convergence rate faster than $\Omega(1/k)$. This result, which stems from the uncertainty over unseen tasks, is formally established in Theorem˜B.1 and serves as a baseline for evaluating the tightness of our upper bounds.
@@ 4.1 Near optimal rates via fixed, horizon-dependent regularization strength
We apply last-iterate convergence results for SGD to the surrogate losses used by IGD under random orderings. Our analysis first considers the fixed-regularization setting, which is common in practice and—unlike our result with increasing regularization—extends cleanly to random orderings without replacement (see Remark˜4.5). Specifically, using the results of Evron et al. [14] together with the smoothness and upper bound from our Lemma˜3.1, we establish the following.
@@
Assume a random with-replacement ordering over jointly realizable tasks. Then, for each of Schemes˜1 and 2, the expected loss after $k\geq 1$ iterations is upper bounded as:
@@
    Fixed coefficient: For Scheme˜1 with a regularization coefficient $\lambda>0$,
@@
    Fixed budget: For Scheme˜2 with step size $\gamma\in(0,1/R^{2})$ and budget $N\in\mathbb{N}$,
@@
All proofs for this subsection are provided in App.˜D.
The rates established in Lemma˜4.3 raise a natural question: What choice of the regularization strength—i.e., the regularization coefficient $\lambda$ or step count NN—achieves the tightest bound?
@@
Assume a random with-replacement ordering over jointly realizable tasks. When the regularization strengths in Lemma˜4.3 are set as follows:
@@
    Fixed coefficient: For Scheme˜1, set regularization coefficient $\lambda\triangleq R^{2}(\ln k-1)$;
@@
    Fixed budget: For Scheme˜2, choose step size $\gamma\in(0,1/R^{2})$ and set budget $N\triangleq\frac{\ln\left({1-\frac{1}{\ln k}}\right)}{\ln\left({1-\gamma R^{2}}\right)}$;
Then, under either Scheme˜1 or Scheme˜2, the expected loss after $k\geq 2$ iterations is bounded as:
@@
The rates in Lemmas˜4.3 and 4.4 extend to random orderings without replacement; see App.˜D for details.
This marks a significant improvement over the $\mathcal{O}(1/k^{1/4})$ rate established by Evron et al. [14] for the unregularized scheme. By tuning the regularization strength, we gain control over the smoothness of the surrogate losses $f^{(t)}_{r}$ and $f^{(t)}_{b}$ in Reductions˜1 and 2, allowing us to attain the $\mathcal{O}(\log k/k)$ rate that is optimal within the SGD framework used in their analysis. In contrast, their unregularized scheme lacked this flexibility, which made achieving such rates considerably more difficult and potentially out of reach. A similar rate can also be derived from the last-iterate bounds of Varre et al. [43], as the smoothness induced by our choice of regularization falls within the applicable regime of their results.
@@ 4.2 Optimal rates via increasing regularization
We present the first result in continual linear regression that achieves the optimal rate for the last iterate, matching the known lower bound. This is obtained by employing a schedule in which the regularization strength increases over time. We discuss these findings and their connections to prior work in Section˜6. All proofs for this subsection are provided in App.˜E.
@@
Assume a random with-replacement ordering over jointly realizable tasks. Consider either Scheme˜1 or Scheme˜2 with the following time-dependent schedules:
@@
    Scheduled coefficient: For Scheme˜1, set regularization coefficient $\lambda_{t}=\frac{13R^{2}}{3}\cdot\frac{k+1}{k-t+2}$;
@@
    For Scheme˜2, choose step sizes $\gamma_{t}\in(0,1/R^{2})$ and set budget $N_{t}=\frac{3}{13\gamma_{t}R^{2}}\cdot\frac{k-t+2}{k+1}$;
Then, under either Scheme˜1 or Scheme˜2, the expected loss after $k\geq 2$ iterations is bounded as:
@@
Establishing the theorem requires a novel last-iterate bound in stochastic optimization, as no existing analysis yields a $\mathcal{O}(1/k)$ guarantee for last-iterate convergence in the realizable setting. A standard path to such rates is to use a decreasing step-size schedule. However, our setting is more nuanced: the quantities we control are the regularization strengths—i.e., the regularization coefficient or step budget in Scheme˜1 or 2—which inherently modify the surrogate objectives $f^{(t)}_{r}$ and $f^{(t)}_{b}$ in Reductions˜1 and 2.
To handle this, we analyze SGD applied to time-varying objectives—a generalization of standard SGD. For this analysis to yield meaningful guarantees, the evolving surrogates must closely approximate the original loss. Indeed, this condition holds, as verified by Lemma˜3.1, thus enabling the application of the next lemma. [^3] While Lemma 4.7 serves Theorem 4.6 in our least-squares setting, it is stated more generally to apply to a broader family of objectives.
@@
Assume $\tau$ is a random with-replacement ordering over $M$ jointly realizable convex and $\beta$-smooth loss functions $f(\cdot;m)\colon\mathbb{R}^{d}\to\mathbb{R}$. Define the average loss $f({\mathbf{w}})\triangleq\mathbb{E}_{m\sim\tau}f({\mathbf{w}};m)$. Let $k\geq 2$, and suppose $\left\{f^{(t)}(\cdot;m)\mid t\in[k],m\in[M]\right\}$ are time-varying surrogate losses that satisfy:
@@
Then, IGD (Scheme˜3) with a diminishing step size that satisfies $\nu_{t}\leq\eta_{t}=\eta\left(\frac{k-t+2}{k+1}\right),~\forall t\in[k]$ for some $\eta\leq 3/(13\beta)$, guarantees the following expected loss bound:
@@ 4.3 Do not forget forgetting: Extension to seen-task loss
We now take the opportunity to briefly revisit our results through the lens of other quantities of interest beyond the average (training) loss defined in Definition˜2.1.
@@
In theoretical work, however, it is essential to define such quantities explicitly. Doan et al. [11] defined forgetting at time $k$ as the drift in model outputs, e.g., $\frac{1}{k}\sum_{t=1}^{k}\left\|{\mathbf{X}_{\tau_{t}}({\mathbf{w}}_{k}-{\mathbf{w}}_{t})}\right\|^{2}$. Nevertheless, this can be large even if the model improves between times $t$ and kk—that is, in the presence of positive backward transfer.
@@
However, in schemes like our regularized approaches (Schemes˜1 and 2), where convergence is not achieved despite realizability, loss degradation can be negative due to backward transfer. As a result, it is sensitive to worst-case analytical “manipulations” and difficult to analyze theoretically.
@@
In App.˜E, we extend Theorem˜4.6 from the average loss to the seen-task loss. Specifically, we show that increasing regularization also achieves an $\mathcal{O}(1/k)$ rate for the expected seen-task loss. But, is this the optimal rate for seen-task loss?
The next lemma shows that, at least under explicit isotropic regularization (Scheme˜1), it is optimal. Proof in App.˜B. More precisely, under random task orderings, no regularization schedule yields a rate faster than $\mathcal{O}(1/k)$ for the expected seen-task loss. In Section˜6, we discuss how non-isotropic regularization—at the cost of additional space complexity—can ensure a seen-task loss of zero.
Lemma 4.9 (Lower bound for seen-task loss of the isotropic Scheme˜1 under any coefficient sequence).
For any $d\geq 2$, initialization ${\mathbf{w}}_{0}\in\mathbb{R}^{d}$, and regularization coefficient sequence $\lambda_{1},\ldots,\lambda_{k}\geq 0$, there exists a set of jointly realizable linear regression tasks $\left\{\left({\mathbf{X}_{m},\mathbf{y}_{m}}\right)\right\}_{m=1}^{M}$ such that, under a with-replacement random task ordering, Scheme˜1 incurs seen-task loss $\mathcal{L}_{1:k}({\mathbf{w}}_{k})=\Omega(1/k)$ with probability at least $1/10$.